# Universal Analog Computation: Fraïssé limits of dynamical systems

### v1

@@ 5. Connections to coalgebra and domain theory
To put things into perspective, we establish connections to the two main areas of research on computation from a general point of view: coalgebra and domain theory. In particular, we make precise the strong analogies to domain theory alluded to in the preceding section.
Coalgebra. The field of coalgebra studies the ‘mathematics of computational dynamics’ [58, vii]. From this perspective, we view our nondeterministic systems $(X,T)$ as the coalgebra $T:X\to\mathsf{F}(X)$ that maps each $x\in X$ to the (closed) set $T(x)$ of its successor states. To make this precise, we need to specify the endofunctor $\mathsf{F}$. Since $X$ is a second-countable Stone space, it is natural to work in that category. So we need to put a topology on the collection $\mathsf{F}(X)$ of all closed subsets of $X$. Such topologies are well-studied as hyperspaces [11]. A standard one is the Vietoris topology: it is generated by (with $U\subseteq X$ open)
@@
Domain theory. [34] studies dynamical systems via domain theory by considering, among others, the upper Vietoris topology on $\mathsf{F}(X)\setminus\{\emptyset\}$ which (in our case where $X$ is a second-countable Stone space) is an $\omega$-algebraic bounded-complete dcpo and $\mathsf{F}(T)$ a Scott-continuous function. [54] provides another way of assigning a dynamical system $(X,T)$ (viewed as a computational process) to a domain $D$ with a continuous $\llbracket T\rrbracket:D\to D$ (viewed as the denotational semantics of that process: a program $\llbracket T\rrbracket$ of type $D$). This approach also incorporates probability measures on $X$. Here, we will make precise the analogy between, on the one hand, the dynamical systems that we study in this paper and, on the other hand, domain theory. To do so, we prove an equivalence between the category $\mathsf{Sys}^{\mathsf{ef}}$} and a ‘domain-theoretic’ category that we call $\mathsf{dynAlg}$.
@@
Here, we are interested in the following subcategory of $\mathsf{Alg}$. Call an element $c$ of an algebraic lattice $A$ a co-atom if $c$ is not the greatest element of $A$, but for all $y\in A$, if $c\leq y$, then either $y=c$ or $y$ is the greatest element of AA. [^6] The order-theoretically dual notion is that of an atom: an element $a\in A$ that is not the least element but for all $y\in A$, if $y\leq a$, then either $y=a$ or $y$ is the least element. For an element $x\in A$, write $\mathrm{CoAt}(x)$ for the set of all co-atoms $c$ in $A$ with $c\geq x$. A function $f:A\to B$ between algebraic lattices is co-atomic if, for all $x\in A$, we have $f(x)=\bigwedge f[\mathrm{CoAt}(x)]$. Now, we define the subcategory $\mathsf{b}\mathsf{Alg}^{\mathsf{c}}$ of $\mathsf{Alg}$ as follows: The objects are the algebraic lattices whose compact elements form a countable sublattice that is a Boolean algebra (i.e., distributive and complemented), and the morphisms are the Scott-continuous and co-atomic functions. [^7] Algebraic lattices whose compact elements form a sublattice are known as arithmetic lattices [41, def. I-4.7]. So we might call the objects of $\mathsf{b}\mathsf{Alg}^{\mathsf{c}}$ ‘Boolean-arithmetic lattices’.
@@
To not digress from establishing the universality results for $\mathsf{Sys}^{\mathsf{ef}}$}, we move the proof to appendix A.2. Of course, the universality result then also holds for $\mathsf{dynAlg}$. Hence the result can be seen within the tradition of existence results concerning universal domains [32], albeit now for the new category of dynamic algebraic lattices. To connect back to coalgebras, note that the objects of $\mathsf{dynAlg}$ are coalgebras for the identity functor on $\mathsf{b}\mathsf{Alg}^{\mathsf{c}}$. However, the morphisms are not coalgebra morphisms. In this case, a coalgebra morphism $\varphi:(A,\alpha)\to(B,\beta)$ would be a morphism $\varphi:A\to B$ in $\mathsf{b}\mathsf{Alg}^{\mathsf{c}}$ such that $\varphi\circ\alpha=\beta\circ\varphi$, yet here we do not have equality, but only the inequality expressed in 5, which, together with 4, forms a weaker equivariance condition. On the other hand, the morphisms in $\mathsf{dynAlg}$ are also stronger in the sense that they are not just a ‘one-directional’ morphism in $\mathsf{b}\mathsf{Alg}^{\mathsf{c}}$ but they rather are ‘bidirectional’: They are embedding-projection pairs, as demanded by 2 and 3, which play an important role in $\mathsf{Alg}$. Thus, this corroborates the analogy to ef-pairs, which we mentioned in section 4.
6. Algebroidal categories of systems
To show our universality result, we will use the Fraïssé theorem (aka Fraïssé–Jónsson theorem) from model theory in its category-theoretic generalization due to [32, thm. 1.1]. (This is further discussed and generalized by [66] and [17].) To do so, we (1) have to show that $\mathsf{Sys}^{\mathsf{ef}}$} is a so-called algebroidal category. Then (2) the universality result reduces to showing that the subcategory of ‘category-theoretically finite objects’ has the so-called joint embedding property and the amalgamation property. In this section, we define all the required terminology and state a sufficient condition for algebroidality, which we show to be satisfied by $\mathsf{Sys}^{\mathsf{ef}}$} (and also for other categories of systems). In the next section, we prove that the condition is indeed sufficient. Then section 8 derives the universality result by establishing the joint embedding property and the amalgamation property.
We recall the definitions (from [32]). Let $\mathsf{C}$ be a category. We write $\omega=\{0,1,2,\ldots\}$. An $\omega$-chain in $\mathsf{C}$ is a structure $(A_{i},f_{i,i+1})_{i\in\omega}$ where each $A_{i}$ is an object in $\mathsf{C}$ and each $f_{i,i+1}:A_{i}\to A_{i+1}$ is a morphism in $\mathsf{C}$ (we will drop the commas in the subscripts). For $i\leq j$ in $\omega$, define $f_{ij}:=f_{j-1j}\circ\ldots\circ f_{ii+1}$, with $f_{ij}:=\mathrm{id}_{A_{i}}$ if $i=j$. So we have the diagram $(A_{i},f_{ij})$ of shape $\omega$ in $\mathsf{C}$. If it has a colimit in $\mathsf{C}$, we denote it $(A,f_{i})$ with $f_{i}:A_{i}\to A$, and we also call it the colimit of the $\omega$-chain $(A_{i},f_{ii+1})$. An object $B$ of $\mathsf{C}$ is (category-theoretically) finite if, for all $\omega$-chains $(A_{i},f_{ii+1})_{i\in\omega}$ with colimit $(A,f_{i})$, if $g:B\to A$ is a morphism, then there exists $n$ such that, for all $i\geq n$, there is a unique morphism $h:B\to A_{i}$ such that $g=f_{i}\circ h$.
A category $\mathsf{C}$ is semi-algebroidal if (1) all its morphisms are monic, (2) every object of $\mathsf{C}$ is a colimit of an $\omega$-chain of finite objects in $\mathsf{C}$, and (3) every $\omega$-chain of finite objects in $\mathsf{C}$ has a colimit in $\mathsf{C}$. Finally, $\mathsf{C}$ is algebroidal if it is semi-algebroidal, it contains up to isomorphism only countable many finite objects, and between any two finite objects there are only countably many morphisms. As useful fact of algebroidal categories is the following.
Lemma 6.1.
Let $\mathsf{C}$ be category in which all morphisms are monic. Let $B$ be a category-theoretically finite object of $\mathsf{C}$. If $(B,f_{i})$ is the colimit of an $\omega$-chain $(A_{i},f_{ii+1})_{i\in\omega}$ in $\mathsf{C}$, then some $f_{i}:A_{i}\to B$ is an isomorphism.
Proof.
@@ 6. Algebroidal categories of systems
Moving to topology, by a countable inverse (or projective) system of topological spaces we mean the data of a countable directed partial order $(I,\leq)$, a family $(X_{i})_{i\in I}$ of topological spaces, and, for each $i\leq j$ in $I$, a surjective map $f_{ij}:X_{j}\to X_{i}$ such that $f_{ii}$ is the identity function on $X_{i}$ and, for $i\leq j\leq k$ in $I$, we have $f_{ij}\circ f_{jk}=f_{ik}$. The category-theoretic limit $(X,f_{i})$, also called inverse or projective limit, is given as
@@
endowed with the subspace topology of the product topology, and $f_{i}\big(\langle x_{i}:i\in I\rangle\big):=x_{i}$ are the canonical projection maps. One can show that the $f_{i}$ are surjective, as is any mediating morphism from another cone. A standard result is (e.g., [22, prop. 3.1]):
@@
We will prove this result in the next section. We end this section by checking to which categories of systems the result applies (theorem 6.4 below) and to which it does not (example 6.5 below).
@@
However, we will see, in section 9, that the category $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$ is semi-algebroidal, but its category-theoretically finite objects are very different: they are the shifts over a finite alphabet. Since there are uncountably many such shifts, $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$ is only semi-algbroidal and not algebroidal. This indicates that the notion of category-theoretical finiteness depends quite subtly on the surrounding category.
7. Proof of the algebroidality theorem
In this section, we prove theorem 6.3 in a sequence of propositions. Regarding part (1) of being algebroidal, we have the following.

### v2

@@ 3.2. Morphisms: Embedding-factor pairs
Remark 3.10 (Simulation and abstraction).
The intuition behind an ef-pair $(e,f):(X,T)\to(Y,S)$ is that the factor $f:Y\to X$ abstracts the ‘detailed’ states $y$ in $Y$ into ‘coarser’ states $x$ in $X$, and the embedding $e:X\to Y$ simulates states $x$ in $X$ by states $e(x)$ in $Y$ (cf. [119]). Several states $y$ (namely those $f^{-1}(x)$) can be identified under the abstraction into a single state $x$. And several states $y$ (namely those in $e(x)$) can simulate the state $x$: if there is a dynamic step $x\xrightarrow{T}x^{\prime}$ in $X$, there are states $y$ and $y^{\prime}$ simulating $x$ and $x^{\prime}$, respectively, and $y\xrightarrow{S}y^{\prime}$.
@@
4. Main results and proof technique
With this framework of dynamical systems in place, we can now state our main results (section 4.1). Then we describe the general proof technique, namely, Fraïssé limits (section 4.2).
4.1. Statement of the main results
To connect to existing theories of computation, section 5 establishes a categorical equivalence between our category $\mathsf{Sys}^{\mathsf{ef}}$} of systems and what we call the category $\mathsf{dynAlg}$ of dynamic algebraic lattices. But our two main results on the existence of universal systems are the following.
To state these results, recall that we are looking for a universal system $(U,T)$, i.e., for any system $(Y,S)$, we need an ef-pair $(e,f):(Y,S)\to(U,T)$. Moreover, ideally the system $(U,T)$ also is ‘homogeneous’ in the intuitive sense that ‘every part of it looks the same’, so it does not have any ‘idiosyncrasies’. To formalize this, recall that a topological space $X$ is homogeneous if, for any two points $x,y\in X$, there is a homeomorphism $h:X\to X$ such that h(x)=yh(x)=y. [^7] Equivalently, the group $G$ of hoemeomorphisms of $X$ acts transitively on $X$. Since points of $X$ are essentially just continuous functions from the one-point space into $X$, we can rephrase this as follows: If $Y$ is the one-point space and $f,g:Y\to X$ are continuous functions, then there is a homeomorphism $h:X\to X$ such that $h\circ g=f$. By demanding this for more spaces $Y$ (e.g., finite spaces), we get a stronger homogeneity. Formally, then, we define these terms as follows.
Definition 4.1.
Let $\mathsf{C}$ be a category and let $\mathsf{D}$ be a full subcategory. An object $U$ in $\mathsf{C}$ is $\mathsf{D}$-universal, if, for any object $A$ in $\mathsf{D}$, there is a morphism $f:A\to U$ (see figure 2(a)). [^8] We follow [39, 138] in this terminology, but [75, 1757] instead speaks of a ‘cofinal’ object, to avoid the category-theoretic connotation that a universal object is canonically determined by a universal property. An object $U$ in $\mathsf{C}$ is $\mathsf{D}$-homogeneous, if, for any object $A$ in $\mathsf{D}$ and morphisms $f,g:A\to U$, there is an isomorphism $h:U\to U$ such that $h\circ g=f$ (see figure 2(b)).
Regarding our existence result for nondeterministic systems, write $\mathsf{Sys}^{\mathsf{ef}}_{\mathsf{fin}}$ for the full subcategory of $\mathsf{Sys}^{\mathsf{ef}}$} consisting of those systems with a finite state space (i.e., those for which we want homogeneity). Then the result says:
Theorem 4.2 (Universal nondeterministic system).
There is an up to isomorphism unique system in $\mathsf{Sys}^{\mathsf{ef}}$} that is $\mathsf{Sys}^{\mathsf{ef}}$}-universal and $\mathsf{Sys}^{\mathsf{ef}}_{\mathsf{fin}}$-homogeneous.
Recall from the end of section 2 that this is a generalization of the result for profinite undirected graphs [64, 20, 47] (more details in remark 7.3 below). Section 7 discusses further properties of the universal nondeterministic system.
Regarding deterministic systems, there cannot be a universal system, as we discuss in section 8. The shifts are the basic building blocks of deterministic systems (i.e., every deterministic system is a limit of shifts). But the obstacle to universality is that there are uncountably many shifts. So this suggests a strategy to build universal systems for large subclasses of deterministic systems. Namely, choose a countable class $S$ of shifts and consider the full subcategory $\mathsf{DS}^{\mathsf{ef}}(S)$ of $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$ consisting of those systems that are limits of shifts in $S$. Recall remark 3.13 that a colimit along ef-pairs is the same thing as a limit along factors. So our formal definition is this:
Definition 4.3.
As mentioned in example 3.2, the most studied classes of shifts are the class $\mathsf{SFT}$ of shifts of finite type and the class $\mathsf{SOF}$ of sofic shifts. So they are natural choices for $S$. Surprisingly, they share a universal system, as our existence result for deterministic systems says:
Theorem 4.4 (Universal deterministic system).
There is a deterministic system $(U,T)$ such that
  1. (1)
    $(U,T)$ is the up to isomorphism unique system in $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SFT})$ that is $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SFT})$-universal and $\mathsf{SFT}$-homogeneous.
  2. (2)
    $(U,T)$ is the up to isomorphism unique system in $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SOF})$ that is $\mathsf{DS}^{\mathsf{ef}}(\mathsf{SOF})$-universal and $\mathsf{SOF}$-homogeneous.
For brevity, we call $(U,T)$ the universal proshift.
An important corollary of a result of [52] is that the universal proshift has the shadowing property (i.e., simulated orbits are close to a true orbit).
[…]
(a) Universal
[…]
(b) Homogeneous
[…]
(c) Joint embedding
[…]
(d) Amalgamation
4.2. Proof technique: Fraïssé limits
To prove our universality results, we will use the Fraïssé theorem (aka Fraïssé–Jónsson theorem) from model theory in its category-theoretic generalization due to [39, thm. 1.1]. (This is further discussed and generalized by [75] and [21].)
We recall the terminology from [39] to state their version of the Fraïssé theorem. Let $\mathsf{C}$ be a category. We write $\omega=\{0,1,2,\ldots\}$. An $\omega$-chain in $\mathsf{C}$ is a structure $(A_{i},f_{i,i+1})_{i\in\omega}$ where each $A_{i}$ is an object in $\mathsf{C}$ and each $f_{i,i+1}:A_{i}\to A_{i+1}$ is a morphism in $\mathsf{C}$ (we will drop the commas in the subscripts). For $i\leq j$ in $\omega$, define $f_{ij}:=f_{j-1j}\circ\ldots\circ f_{ii+1}$, with $f_{ij}:=\mathrm{id}_{A_{i}}$ if $i=j$. So we have the diagram $(A_{i},f_{ij})$ of shape $\omega$ in $\mathsf{C}$. If it has a colimit in $\mathsf{C}$, we denote it $(A,f_{i})$ with $f_{i}:A_{i}\to A$, and we also call it the colimit of the $\omega$-chain $(A_{i},f_{ii+1})$. An object $B$ of $\mathsf{C}$ is (category-theoretically) finite if, for all $\omega$-chains $(A_{i},f_{ii+1})_{i\in\omega}$ with colimit $(A,f_{i})$, if $g:B\to A$ is a morphism, then there exists $n$ such that, for all $i\geq n$, there is a unique morphism $h:B\to A_{i}$ such that $g=f_{i}\circ h$.
Definition 4.5.
A category $\mathsf{C}$ is semi-algebroidal if
  1. (1)
    all morphisms in $\mathsf{C}$ are monic,
  2. (2)
    every object of $\mathsf{C}$ is a colimit of an $\omega$-chain of finite objects in $\mathsf{C}$, and
  3. (3)
    every $\omega$-chain of finite objects in $\mathsf{C}$ has a colimit in $\mathsf{C}$.
Finally, $\mathsf{C}$ is algebroidal if it is semi-algebroidal, it contains up to isomorphism only countably many finite objects, and between any two finite objects there are only countably many morphisms.
The term ‘algebroidal’ traces back, as [39, 138] note, to [109] (in domain theory) and [8] (in category theory). Indeed, [109, 780] see algebroidal categories as a generalization of $\omega$-algebraic cpo’s, which are a common kind of partial order studied in domain theory. [^9] We will not discuss cpo’s later, but we give here a definition for completeness and to highlight that it is a special case of an algebroidal category. A cpo is a partial order $(P,\leq)$ with a least element in which every increasing $\omega$-chain has a least upper bound (i.e., ‘colimit’). An element $b\in P$ is $\omega$-compact (i.e., category-theoretically finite) if any increasing $\omega$-chain $(a_{n})$ whose least upper bound is $\geq b$ contains an element $a_{n}$ which is $\geq b$. Finally, $P$ is $\omega$-algebraic if it contains only countably many compact elements, and for any element $a\in P$, there is an increasing $\omega$-chain of compact elements whose least upper bound is $a$. In section 5, we give domain-theoretic definitions in (the more common) terms of directed subsets rather than $\omega$-chains, but these are equivalent if there are only countably many compact elements [1, sec. 2.2.4]. Thus, $\omega$-algebraic cpo’s—when viewed as a poset category [^10] The objects are the elements of the partial order $(P,\leq)$ and there is exactly one morphism from $a$ to $b$ if $a\leq b$, with no further morphisms.—provide a simple example of algebroidal categories. Other common examples also arise in domain theory: many categories of domains form algebroidal categories [39].
Definition 4.6.
Let $\mathsf{C}$ be a category. Then $\mathsf{C}$ has the joint embedding property if for any objects $A$ and $B$, there is an object $C$ with morphisms $f:A\to C$ and $g:B\to C$ (see figure 2(c)). And $\mathsf{C}$ has the amalgamation property if for any objects $A$, $B$, and $B^{\prime}$ with morphisms $f:A\to B$ and $f^{\prime}:A\to B^{\prime}$, there is an object $C$ with morphisms $g:B\to C$ and $g^{\prime}:B^{\prime}\to C$ such that $g\circ f=g^{\prime}\circ f^{\prime}$ (see figure 2(d)).
Now the category-theoretic Fraïssé theorem says:
Theorem 4.7 ([39]).
Thus, our proof strategy, both for nondeterministic and deterministic systems, will be to show that the relevant category of systems (1) is algebroidal and (2) that the full subcategory of category-theoretically finite systems has the joint embedding property and the amalgamation property. We will achieve (1) by proving ‘algebroidality theorems’, i.e., theorems providing easily check-able sufficient conditions for a category of systems to be (semi-) algebroidal (theorems 6.2, 8.6, and 9.2).
@@ 5. Connections to coalgebra and domain theory
To put things into perspective, we establish connections to the two main areas of research on computation from a general point of view: coalgebra and domain theory. In particular, we make precise the strong analogies to domain theory alluded to in section 3.
Coalgebra. The field of coalgebra studies the ‘mathematics of computational dynamics’ [65, vii]. From this perspective, we view our nondeterministic systems $(X,T)$ as the coalgebra $T:X\to\mathsf{F}(X)$ that maps each $x\in X$ to the (closed) set $T(x)$ of its successor states. To make this precise, we need to specify the endofunctor $\mathsf{F}$. Since $X$ is a second-countable Stone space, it is natural to work in that category. So we need to put a topology on the collection $\mathsf{F}(X)$ of all closed subsets of $X$. Such topologies are well-studied as hyperspaces [14]. A standard one is the Vietoris topology: it is generated by (with $U\subseteq X$ open)
@@
Domain theory. [41] studies dynamical systems via domain theory by considering, among others, the upper Vietoris topology on $\mathsf{F}(X)\setminus\{\emptyset\}$ which (in our case where $X$ is a second-countable Stone space) is an algebraic domain and $\mathsf{F}(T)$ a Scott-continuous function (definitions below). [^11] Moreover, the algebraic domain has only countably many compact elements (definition below) and is bounded-complete (any two elements with an upper bound have a least upper bound). [61] provides another way of assigning a dynamical system $(X,T)$ (viewed as a computational process) to a domain $D$ with a continuous $\llbracket T\rrbracket:D\to D$ (viewed as the denotational semantics of that process: a program $\llbracket T\rrbracket$ of type $D$). This approach also incorporates probability measures on $X$. Here, we will make precise the analogy between, on the one hand, the dynamical systems that we study in this paper and, on the other hand, domain theory. To do so, we prove an equivalence between the category $\mathsf{Sys}^{\mathsf{ef}}$} and a ‘domain-theoretic’ category that we call $\mathsf{dynAlg}$.
@@
Here, we are interested in the following subcategory of $\mathsf{Alg}$. Call an element $c$ of an algebraic lattice $A$ a co-atom if $c$ is not the greatest element of $A$, but for all $y\in A$, if $c\leq y$, then either $y=c$ or $y$ is the greatest element of AA. [^13] The order-theoretically dual notion is that of an atom: an element $a\in A$ that is not the least element but for all $y\in A$, if $y\leq a$, then either $y=a$ or $y$ is the least element. For an element $x\in A$, write $\mathrm{CoAt}(x)$ for the set of all co-atoms $c$ in $A$ with $c\geq x$. A function $f:A\to B$ between algebraic lattices is co-atomic if, for all $x\in A$, we have $f(x)=\bigwedge f[\mathrm{CoAt}(x)]$. Now, we define the subcategory $\mathsf{b}\mathsf{Alg}^{\mathsf{c}}$ of $\mathsf{Alg}$ as follows: The objects are the algebraic lattices whose compact elements form a countable sublattice that is a Boolean algebra (i.e., distributive and complemented), and the morphisms are the Scott-continuous and co-atomic functions. [^14] Algebraic lattices whose compact elements form a sublattice are known as arithmetic lattices [48, def. I-4.7]. So we might call the objects of $\mathsf{b}\mathsf{Alg}^{\mathsf{c}}$ ‘Boolean-arithmetic lattices’.
@@
To not digress from establishing the universality results for $\mathsf{Sys}^{\mathsf{ef}}$}, we move the proof to appendix A.2. Of course, the universality result then also holds for $\mathsf{dynAlg}$. Hence the result can be seen within the tradition of existence results concerning universal domains [39], albeit now for the new category of dynamic algebraic lattices. To connect back to coalgebras, note that the objects of $\mathsf{dynAlg}$ are coalgebras for the identity functor on $\mathsf{b}\mathsf{Alg}^{\mathsf{c}}$. However, the morphisms are not coalgebra morphisms. In this case, a coalgebra morphism $\varphi:(A,\alpha)\to(B,\beta)$ would be a morphism $\varphi:A\to B$ in $\mathsf{b}\mathsf{Alg}^{\mathsf{c}}$ such that $\varphi\circ\alpha=\beta\circ\varphi$, yet here we do not have equality, but only the inequality expressed in 5, which, together with 4, forms a weaker equivariance condition. On the other hand, the morphisms in $\mathsf{dynAlg}$ are also stronger in the sense that they are not just a ‘one-directional’ morphism in $\mathsf{b}\mathsf{Alg}^{\mathsf{c}}$ but they rather are ‘bidirectional’: They are embedding-projection pairs, as demanded by 2 and 3, which play an important role in $\mathsf{Alg}$. Thus, this corroborates the analogy to ef-pairs, which we mentioned in section 3.2.
6. Algebroidal categories of nondeterministic systems
We now work toward the universality result for nondeterministic systems. As described in the proof strategy (section 4.2), we start with an algebroidality theorem: sufficient conditions for when a category of (nondeterministic) systems is algebroidal. We state the theorem in section 6.1 and prove it in section 6.2. In the next section, we use this to prove the existence of a universal nondeterministic system.
6.1. Algebroidality theorem for nondeterministic systems
@@
To state the theorem, let us first recall some terminology from topology. By a countable inverse (or projective) system of topological spaces we mean the data of a countable directed partial order $(I,\leq)$, a family $(X_{i})_{i\in I}$ of topological spaces, and, for each $i\leq j$ in $I$, a surjective map $f_{ij}:X_{j}\to X_{i}$ such that $f_{ii}$ is the identity function on $X_{i}$ and, for $i\leq j\leq k$ in $I$, we have $f_{ij}\circ f_{jk}=f_{ik}$. The category-theoretic limit $(X,f_{i})$, also called inverse or projective limit, is given as
@@
endowed with the subspace topology of the product topology, and $f_{i}\big(\langle x_{i}:i\in I\rangle\big):=x_{i}$ are the canonical projection maps. One can show that the $f_{i}$ are surjective, as is any mediating morphism from another cone. A standard result is (e.g., [28, prop. 3.1]):
@@
We will prove this result in section 6.2. We end this subsection by checking to which categories of systems the result applies (theorem 6.3 below) and to which it does not (example 6.4 below).
@@
However, we will see, in section 8, that the category $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$ is semi-algebroidal, but its category-theoretically finite objects are very different: they are the shifts. Since there are uncountably many such shifts, $\mathsf{det}\mathsf{Sys}^{\mathsf{ef}}$ is only semi-algebroidal and not algebroidal. This indicates that the notion of category-theoretical finiteness depends quite subtly on the surrounding category.
6.2. Proof of the algebroidality theorem
In this section, we prove theorem 6.2 in a sequence of propositions. (We will use some of these again when considering deterministic systems.)
Regarding part (1) of being algebroidal, we have the following.